OpenStudy (keenan):

absolute value of 3x-8 greaterthan or equal to 5 is........ two solutions

OpenStudy (anonymous):

Just remember your absolute value rules.. \[\forall a,b \in \mathbb{R} \text{ and } b > 0,\ \ |a| > b \implies a > b \text{ or } a < -b\] So in this case you have: \[|3x -8| \ge 5\] Which part of this relation is your a, and which is your b?

OpenStudy (keenan):

its dealing with inequalities

OpenStudy (anonymous):

I've written inequalities..

OpenStudy (anonymous):

Compare this: \[|a| \ge b\] To this: \[|3x -8| \ge 5\] Which part in your problem corresponds to the 'a' part and which corresponds to the 'b'?

OpenStudy (keenan):

a=3x-8, b=5 c= the value of x

OpenStudy (anonymous):

There is no c..

OpenStudy (keenan):

then what is the value of x

OpenStudy (anonymous):

Ok, so if a is 3x - 8 and b is 5, and \[|a| \ge b \implies a \ge b \text{ or } a \le -b\] So can you write your two new inequalities using that formula and your a and b?

OpenStudy (keenan):

ok but how

OpenStudy (anonymous):

I just told you.

OpenStudy (anonymous):

\[|a| \ge b\] Implies two inequalities: \[a \ge b\] or \[a \le -b\]

OpenStudy (keenan):

oh ok nevermind i got it

OpenStudy (keenan):

thank you

OpenStudy (anonymous):

So what would your two inequalites be?